OpenStudy (anonymous):

Evaluate m – n when m = –2.6 and n = 7.9. A. –10.5 B. –5.3 C. 5.3 D. 10.5

OpenStudy (anonymous):

i will fan and medal

OpenStudy (anonymous):

@kourtney2014

geerky42 (geerky42):

Just plug in values. \(\color{blue}m - \color{red}n\quad\longrightarrow\quad\color{blue}{-2.6}-\color{red}{7.9}=?\)

OpenStudy (anonymous):

i dont have a calculator

OpenStudy (anonymous):

b

OpenStudy (anonymous):

-5.3

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

wait no d

OpenStudy (anonymous):

is it b or d??

OpenStudy (anonymous):

no b i mean

OpenStudy (anonymous):

ok, its -10.5 so d its because it has a negative and with a negative and positive it equals a negative. your changing the minus sign and putting a positive sign so -2.6+7.9=-10.5

OpenStudy (anonymous):

i put b and it was right

OpenStudy (anonymous):

ok. im sorry it was just that the negatives always get to me and yea.

OpenStudy (anonymous):

can you help me one more time please?????????????

OpenStudy (anonymous):

Line BC contains points B (4, −5) and C (3, 2). Line DE contains points D (2, 0) and E (9, 1). Lines BC and DE are Select one: a. parallel b. perpendicular c. neither

OpenStudy (anonymous):

its fine and yea

geerky42 (geerky42):

So it's -5.3 and not -10.5? lol

OpenStudy (anonymous):

ok. i just posted teh question

OpenStudy (anonymous):

and yes @geerky42

geerky42 (geerky42):

lol okay...

OpenStudy (anonymous):

its parral @kourtney2014

geerky42 (geerky42):

How do you know?

OpenStudy (anonymous):

ok thanks @pizza4dayz

OpenStudy (anonymous):

The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5) a. y = 5x − 15 b. y = −5x + 15 c. y = x + d. y = x −